1. Find the slope: y=-3/2x + 5/3
2. Find the equation: slope-4 through point (-2,3)

Answer by checkley71(8403)

1) Y=-3X/2=5/3 THE LINE FORMULA IS [Y=mX+b] THUS THE SLOPE (m)=-3/2
2) 3=-2*-4+b OR 3=8+b OR b=3-8 OR b=-5 THUS THE EQUATION IS [Y=-4X-5]